In a movie scene, a car was driven at a speed of 30 m s-1

up a ramp inclined at 45o
to the ground. The car ran off the top of the ramp which is 2.0 m above the ground.
Assuming that air resistance is negligible, calculate
a) the time the car is in the air.
b) the horizontal distance-travelled by the car from the ramp to the point of
impact on the ground.
1
Expert's answer
2021-08-16T14:44:13-0400

Given:

"v_{0x}=v_0\cos\theta"

"v_{0y}=v_0\sin\theta"

"v_0=30\:\rm m/s"

"y_0=2.0\:\rm m"

"\theta=45^{\circ}"


The equations of motion of a car

"x(t)=v_{0x}*t=30\cos45^{\circ}*t
=21t""y(t)=y_0+v_{0y}*t-\frac{gt^2}{2}=2.0+30\sin45^{\circ}-\frac{9.8*t^2}{2}\\
=2.0+21t-4.9t^2"

(a)

"2.0+21t-4.9t^2=0"

"t=4.4\:\rm s"

(b)

"L=x(4.4\:\rm s)=21*4.4=93\:\rm m"